# SIC-POVMs and orders of real quadratic fields

### v1

@@ 5.2. Statement of the Geometric Count Conjecture
The following conjecture counts the number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$.
@@ 5.3. Verification of cases of the Geometric Count Conjecture
We have verified that the two sets in the Geometric Count conjecture have the same size for dimensions $d\leq 90$, $d\neq 3$ under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39]. Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are Weyl–Heisenberg SIC fiducials. By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs.
The dimensions $d\in\{1,2\}$ are special cases—the only dimensions for which $\Delta<0$, or equivalently, the conjugacy classes in $\mathcal{C}_{d-1}$ are elliptic. For $d=1$, there is exactly one SIC ($\mathbb{C}^{1}$ as a line in $\mathbb{C}^{1}$), and it is (trivially) Weyl–Heisenberg covariant. The associated order $\mathcal{O}_{-4}=\mathbb{Z}[i]$ is a maximal order of class number one, so the conjecture is true. For $d=2$, there is again a unique SIC up to unitary transformations (a regular tetrahedron on the Riemann sphere $\mathbb{P}^{1}(\mathbb{C})$), and it may be chosen to be Weyl–Heisenberg covariant. The associated order $\mathcal{O}_{-3}=\mathbb{Z}[e^{2\pi i/3}]$ is a maximal order of class number one, so the conjecture is true.
5.4 omits the case $d=3$, where $\Delta=0$, where associated field $K_{\!\Delta}=\mathbb{Q}$ has $\mathcal{O}_{\!\Delta}=\operatorname{Z}$. In this case, there is a continuous one-parameter family of Weyl–Heisenberg covariant SICs [39]. For $d=3$ the classification (3) above is the set $\mathcal{C}_{2}$ of parabolic conjugacy classes of $\mathbf{SL}_{2}(\mathbb{Z})$, which is infinite.
@@
For $4\leq d\leq 90$ we computed $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ using Magma [17] by computing the equivalent quantity $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$. Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$. For each $f^{\prime}\mid f$, we computed the set of reduced primitive binary quadratic forms of discriminant $(f^{\prime})^{2}\Delta_{0}$ using the Magma function ReducedForms(), written by David Kohel. This function produces a list of $\mathbf{SL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms in which a twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class is always represented either by a single form $ax^{2}+bxy+cy^{2}$ with $a>0$ or a pair of forms $\{ax^{2}+bxy+cy^{2},-ax^{2}+bxy-cy^{2}\}$. Removing forms with $a<0$ produced a list containing one representative for each primitive twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class. Multiplying by $f/f^{\prime}$ and joining the lists gave a list of representatives of all (not necessarily primitive) classes in $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$.
We observe that for each $d$ ($4\leq d\leq 90$) that the size of
@@
$\operatorname{Clm}(\mathcal{O}_{\Delta})=\sum_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (5.5)
is equal to the sum of the corresponding values in the “#” column of Scott’s table beginning on page 3 of [38], which give the number of elements of ${\rm WHSIC}(d)/\operatorname{PEC}(d)$. This is agreement is shown in Table 1. The table also includes the sizes of the ring class groups $\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (equivalently, the number of twisted $\mathbf{GL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms with discriminant $(f^{\prime})^{2}\Delta_{0}$).
$d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\##
$1$ $-4$ $1$ $1$ $1$
$2$ $-3$ $1$ $1$ $1$
$3$ 00 $\ast$ $\ast$ $\infty$
$4$ $5$ $1$ $1$ $1$
$5$ $12$ $1$ $1$ $1$
$6$ $21$ $1$ $1$ $1$
$7$ $8$ $1,2$ $1,1$ $2$
$8$ $5$ $1,3$ $1,1$ $2$
$9$ $60$ $1$ $2$ $2$
$10$ $77$ $1$ $1$ $1$
$11$ $24$ $1,2$ $1,2$ $3$
$12$ $31$,3131,31 , 3 $1,1$ $2$
$13$ $140$ $1$ $2$ $2$
$14$ $165$ $1$ $2$ $2$
$15$ $21$,2,41241,2,41 , 2 , 4 $1,1,2$ $4$
$16$ $221$ $1$ $2$ $2$
$17$ $28$ $1,3$ $1,2$ $3$
$18$ $285$ $1$ $2$ $2$
$19$ $5$ $1,2,4,8$ $1,1,1,2$ $5$
$20$ $357$ $1$ $2$ $2$
$21$ $44$ $1,3$ $1,4$ $5$
$22$ $437$ $1$ $1$ $1$
$23$ $201$,2121,21 , 2 $2,4$ $6$
$24$ $21$ $1,5$ $1,2$ $3$
$25$ $572$ $1$ $2$ $2$
$26$ $69$ $1,3$ $1,3$ $4$
$27$ $681$,2121,21 , 2 $2,4$ $6$
$28$ $29$ $1,5$ $1,2$ $3$
$29$ $780$ $1$ $4$ $4$
$30$ $93$ $1,3$ $1,3$ $4$
$31$ $56$ $1,2,4$ $1,2,4$ $7$
$32$ $957$ $1$ $2$ $2$
$34$ $1020$ $1$ $4$ $4$
$34$ $1085$ $1$ $2$ $2$
$35$ $8$ $1,2,3,4,6,12$ $1,1,1,1,2,4$ $10$
$36$ $1221$ $1$ $4$ $4$
$37$ $1292$ $1$ $4$ $4$
$38$ $1365$ $1$ $4$ $4$
$39$ $40$ $1,2,3,6$ $2,2,2,4$ $10$
$40$ $1517$ $1$ $2$ $2$
$41$ $1596$ $1$ $8$ $8$
$42$ $1677$ $1$ $4$ $4$
$43$ $440$ $1,2$ $2,4$ $6$
$44$ $205$ $1,3$ $2,4$ $6$
$45$ $1932$ $1$ $4$ $4$
 $d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\## $46$ $2021$ $1$ $3$ $3$ $47$ $33$ $1,2,4,8$ $1,1,2,4$ $8$ $48$ $5$ $1,3,7,21$ $1,1,1,4$ $7$ $49$ $92$ $1,5$ $1,6$ $7$ $50$ $2397$ $1$ $2$ $2$ $51$ $561$,2,41241,2,41 , 2 , 4 $2,4,8$ $14$ $52$ $53$ $1,7$ $1,3$ $4$ $53$ $21$,3,$5,1$ 3 , 5 , 15 $1,1,2,6$ $10$ $54$ $2805$ $1$ $4$ $4$ $55$ $728$ $1,2$ $2,4$ $6$ $56$ $3021$ $1$ $6$ $6$ $57$ $348$ $1,3$ $2,6$ $8$ $58$ $3245$ $1$ $4$ $4$ $59$ $840$ $1,2$ $4,8$ $12$ $60$ $3477$ $1$ $4$ $4$ $61$ $3596$ $1$ $6$ $6$ $62$ $413$ $1,3$ $1,4$ $5$ $63$ $60$ $1,2,4,8$ $2,2,4,8$ $16$ $64$ $3965$ $1$ $4$ $4$ $65$ $4092$ $1$ $8$ $8$ $66$ $469$ $1,3$ $3,6$ $9$ $67$ $71$,2,4,8,161248161,2,4,8,161 , 2 , 4 , 8 , 16 $1,1,1,2,4$ $9$ $68$ $4485$ $1$ $4$ $4$ $69$ $4620$ $1$ $8$ $8$ $70$ $4757$ $1$ $5$ $5$ $71$ $361$,2,3,612361,2,3,61 , 2 , 3 , 6 $2,4,4,8$ $18$ $72$ $5037$ $1$ $4$ $4$ $73$ $5180$ $1$ $4$ $4$ $74$ $213$ $1,5$ $1,6$ $7$ $75$ $521$,2,3,612361,2,3,61 , 2 , 3 , 6 $1,2,4,8$ $15$ $76$ $5621$ $1$ $6$ $6$ $77$ $5772$ $1$ $8$ $8$ $78$ $237$ $1,5$ $1,6$ $7$ $79$ $380$ $1,2,4$ $2,4,8$ $14$ $80$ $77$ $1,3,9$ $1,2,6$ $9$ $81$ $6396$ $1$ $12$ $12$ $82$ $6557$ $1$ $3$ $3$ $83$ $051$,2,4,812481,2,4,81 , 2 , 4 , 8 $2,2,4,8$ $16$ $84$ $85$ $1,3,9$ $2,2,6$ $10$ $85$ $7052$ $1$ $4$ $4$ $86$ $7221$ $1$ $10$ $10$ $87$ $8481$,2121,21 , 2 $4,8$ $12$ $88$ $7565$ $1$ $4$ $4$ $89$ $860$ $1,3$ $2,8$ $10$ $90$ $7917$ $1$ $4$ $4$

### v3

@@ 5.2. Statement of the Geometric Count Conjecture
The following conjecture counts the number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$. It is implied by 1.2.
@@
5.4 immediately gives an upper bound for the total number of Weyl–Heisenberg line-SICs in dimensions $d\geq 4$.
Proposition 5.5.
5.4 implies that the number $\left|\operatorname{WHSIC}(d)\right|$ of Weyl–Heisenberg line-SICs in dimensions $d\geq 4$ satisfies
$\left|\operatorname{WHSIC}(d)\right|\leq\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|.$ (5.6)
Proof.
The number of distinct Weyl–Heisenberg line-SICs in a $\operatorname{PEC}(d)$-equivalence class $[S]$ is $\frac{\left|\operatorname{PEC}(d)\right|}{\left|\operatorname{stab}(S)\right|}$, where $\operatorname{stab}(S)$ is the subgroup of $\operatorname{PEC}(d)$ stabilizing a line-SIC $S\in[S]$. The cardinality $\left|\operatorname{stab}(S)\right|$ is independent of the choice of $S\in[S]$. The Weyl–Heisenberg group $\operatorname{WH}(d)$ acts transitively on the set of $d^{2}$ complex lines in the $\operatorname{WH}(d)$-orbit of a fiducial vector, so $\operatorname{WH}(d)$ is a subgroup of $\operatorname{stab}(S)$. Thus $d^{2}$ divides $\left|\operatorname{stab}(S)\right|$. Hence we have $\frac{\left|\operatorname{PEC}(d)\right|}{\left|\operatorname{stab}(S)\right|}\leq\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}$. The result now follows from 5.4. ∎
Remark 5.6.
Table 5.1 presents, in the column labeled $\left|\operatorname{WHSIC}(d)\right|$, empirical data for the number of observed (numerical) SICs for $4\leq d\leq 50$ from Scott and Grassl [41, Table 1]. That table also observes that all known numerical Weyl–Heisenberg SICs of dimensions $4\leq d\leq 50$ have an additional stabilizer element of order $3$. If this statement were true in general, then we could strengthen 5.5 to conclude
$\left|\operatorname{WHSIC}(d)\right|\leq\frac{1}{3}\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$ (5.7)
holds for all $d\geq 4$. Examination of Table 5.1 shows for known numerical SICs that (5.7) holds for $4\leq d\leq 50$ and that equality in (5.7) holds for
$d\in\{5,13,17,23,25,27,29,37,41,43,47,49\}.$ (5.8)
@@ 5.3. Verification of cases of the Geometric Count Conjecture
We have verified that the two sets in the Geometric Count conjecture have the same size for dimensions $4\leq d\leq 90$, under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions, given in Scott [40], are complete. (The list builds on earlier work of Scott and Grassl [41].)
Proposition 5.7 (Empirical Proposition 1).
For dimensions $d$ with $\Delta=(d+1)(d-3)$, the total number of $\operatorname{PEC}(d)$ equivalence classes $[S]$ of numerical SICs equals $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ for $4\leq d\leq 50$ in the Scott and Grassl list [41], and for $51\leq d\leq 90$ in the Scott list [40].
Proof.
For $4\leq d\leq 90$, we computed $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ using Magma [16] by computing the equivalent quantity $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$.
Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$. Recall that for any quadratic order
$\operatorname{Clm}(\mathcal{O}_{\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}).$ (5.9)
@@
For each $f^{\prime}\mid f$, we computed the set of reduced primitive binary quadratic forms of discriminant $(f^{\prime})^{2}\Delta_{0}$ using the Magma function ReducedForms(), written by David Kohel. This function produces a list of $\mathbf{SL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms in which a twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class is always represented either by a single form $ax^{2}+bxy+cy^{2}$ with $a>0$ or a pair of forms $\{ax^{2}+bxy+cy^{2},-ax^{2}+bxy-cy^{2}\}$. Removing forms with $a<0$ produced a list containing one representative for each primitive twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class. Multiplying by $f/f^{\prime}$ and joining the lists gave a list of representatives of all (not necessarily primitive) classes in $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$. The counts of quantities for the different $f^{\prime}$ are given are the unordered set of all ring class numbers $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$. They are given in column labeled by $h_{(f^{\prime})^{2}\Delta_{0}}$ in Table 5.1. The total counts of class numbers are , are given in the column labeled by $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$.
Numerical Weyl–Heisenberg SICs are systems of $d^{2}$ complex lines that are the Weyl–Heisenberg orbit of a single numerical fiducial vector and that appears to have equal angles to $10$ decimal places. Numerical SICs are produced as output of an optimization algorithm. Empirical counts for the number of $\operatorname{PEC}(d)$-equivalence classes of numerical SICs were given by Scott and Grassl for $4\leq d\leq 50$ [41, Tab. 1 and Tab. 2] and by Scott [40, pp. 3–6] for $51\leq d\leq 90$. The classes in the tables are denoted by dimension and letter, eg., $31a$. Tab. 1 of [41] contains the subset of exact SICs, grouped into Galois multiplets, while Tab. 2 of [41] gives a list of numerical SICs, not grouped into Galois multiplets. In both cases one can count the total number of $\operatorname{PEC}(d)$-equivalence classes.
We also found agreement for all dimensions $51\leq d\leq 90$ of $\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$ given in Table 5.2 with the total count of $\operatorname{PEC}(d)$-classes of SICs in each dimension listed in [40, Appendix C]. ∎
@@
$d$$\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}h_{(f^{\prime})^{2}\Delta_{0}}$ $\Delta$ $\Delta_{0}$ $f^{\prime}$ $\sigma_{0}(f)$ $h_{(f^{\prime})^{2}\Delta_{0}}$ $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$
$1$ $-4$ $-4$ $1$ $1$ $1$ $1$
$2$ $-3$ $-3$ $1$ $1$ $1$ $1$
$3$ 00 00 $\mathbb{N}$infty $1$infty
$4$ $5$ $5$ $1$ $1$ $1$ $1$
$5$ $12$ $12$ $1$ $1$ $1$ $1$
$6$ $21$ $21$ $1$ $1$ $1$ $1$
$7$ $32$ $8$ $1,2$ $2$ $1,1$ $2$
$8$ $45$ $5$ $1,3$ $2$ $1,1$ $2$
$9$ $60$ $60$ $1$ $1$ $2$ $2$
$10$ $77$ $77$ $1$ $1$ $1$ $1$
$11$ $96$ $24$ $1,2$ $2$ $1,2$ $3$
$12$ $117$ $31$,3131,31 , 3 $2$ $1,1$ $2$
$13$ $140$ $140$ $1$ $1$ $2$ $2$
$14$ $165$ $165$ $1$ $1$ $2$ $2$
$15$ $192$ $21$,2,41241,2,41 , 2 , 4 $3$ $1,1,2$ $4$
$16$ $221$ $221$ $1$ $1$ $2$ $2$
$17$ $252$ $28$ $1,3$ $2$ $1,2$ $3$
$18$ $285$ $285$ $1$ $1$ $2$ $2$
$19$ $320$ $5$ $1,2,4,8$ $4$ $1,1,1,2$ $5$
$20$ $357$ $357$ $1$ $1$ $2$ $2$
$21$ $396$ $44$ $1,3$ $2$ $1,4$ $5$
$22$ $437$ $437$ $1$ $1$ $1$ $1$
$23$ $480$ $201$,2121,21 , 2 $2$ $2,4$ $6$
$24$ $525$ $21$ $1,5$ $2$ $1,2$ $3$
$25$ $572$ $572$ $1$ $1$ $2$ $2$
$26$ $621$ $69$ $1,3$ $2$ $1,3$ $4$
$27$ $672$ $681$,2121,21 , 2 $2$ $2,4$ $6$
$28$ $725$ $29$ $1,5$ $2$ $1,2$ $3$
$29$ $780$ $780$ $1$ $1$ $4$ $4$
$30$ $837$ $93$ $1,3$ $2$ $1,3$ $4$
$31$ $896$ $56$ $1,2,4$ $3$ $1,2,4$ $7$
$32$ $957$ $957$ $1$ $1$ $2$ $2$
$34$ $1020$ $1020$ $1$ $1$ $4$ $4$
$34$ $1085$ $1085$ $1$ $1$ $2$ $2$
$35$ $1152$ $8$ $1,2,3,4,6,12$ $6$ $1,1,1,1,2,4$ $10$
$36$ $1221$ $1221$ $1$ $1$ $4$ $4$
$37$ $1292$ $1292$ $1$ $1$ $4$ $4$
$38$ $1365$ $1365$ $1$ $1$ $4$ $4$
$39$ $1440$ $40$ $1,2,3,6$ $4$ $2,2,2,4$ $10$
$40$ $1517$ $1517$ $1$ $1$ $2$ $2$
$41$ $1596$ $1596$ $1$ $1$ $8$ $8$
$42$ $1677$ $1677$ $1$ $1$ $4$ $4$
$43$ $1760$ $440$ $1,2$ $2$ $2,4$ $6$
$44$ $1845$ $205$ $1,3$ $2$ $2,4$ $6$
$45$ $1932$ $1932$ $1$ $1$ $4$ $4$
$46$ $2021$ $2021$ $1$ $1$ $3$ $3$
$47$ $2112$ $33$ $1,2,4,8$ $4$ $1,1,2,4$ $8$
$48$ $2205$ $5$ $1,3,7,21$ $4$ $1,1,1,4$ $7$
$49$ $2300$ $92$ $1,5$ $2$ $1,6$ $7$
$50$ $2397$ $2397$ $1$ $1$ $2$ $2$
 $d$$\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}h_{(f^{\prime})^{2}\Delta_{0}}$ $\left|\operatorname{WHSIC}(d)\right|$ $\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}$#\## $1$ $1$ $1$ $1$ $2$ $2$ $12$ $1$ $3$infty $48$infty $4$ $16$ $96$ $1$ $5$ $80$ $240$ $1$ $6$ $96$ $288$ $1$ $7$ $336$ $672$ $2$ $8$ $320$ $768$ $2$ $9$ $864$ $1296$ $2$ $10$ $480$ $1440$ $1$ $11$ $2640$ $2640$ $3$ $12$ $1152$ $2304$ $2$ $13$ $2912$ $4368$ $2$ $14$ $2688$ $4032$ $2$ $15$ $6720$ $5760$ $4$ $16$ $4096$ $6144$ $2$ $17$ $9792$ $9792$ $3$ $18$ $5184$ $7776$ $2$ $19$ $16720$ $13680$ $5$ $20$ $7680$ $11520$ $2$ $21$ $26880$ $16128$ $5$ $22$ $5280$ $15840$ $1$ $23$ $48576$ $24288$ $6$ $24$ $15360$ $18432$ $3$ $25$ $20000$ $30000$ $2$ $26$ $34944$ $26208$ $4$ $27$ $69984$ $34992$ $6$ $28$ $26880$ $32256$ $3$ $29$ $64960$ $48720$ $4$ $30$ $46080$ $34560$ $4$ $31$ $138880$ $59520$ $7$ $32$ $32768$ $49152$ $2$ $33$ $84480$ $63360$ $4$ $34$ $39168$ $58752$ $2$ $35$ $235200$ $80640$ $10$ $36$ $82944$ $62208$ $4$ $37$ $134976$ $101232$ $4$ $38$ $109440$ $82080$ $4$ $39$ $314496$ $104832$ $10$ $40$ $61440$ $92160$ $2$ $41$ $367360$ $137360$ $8$ $42$ $129024$ $96768$ $4$ $43$ $317856$ $158928$ $6$ $44$ $253440$ $126720$ $6$ $45$ $207360$ $155520$ $4$ $46$ $145728$ $145728$ $3$ $47$ $553472$ $207552$ $8$ $48$ $276480$ $147456$ $7$ $49$ $537824$ $230496$ $7$ $50$ $120000$ $180000$ $2$